OpenStudy (anonymous):

3x^2-14x+8, is this factorable?

OpenStudy (cj49):

\[-b \pm \sqrt{b ^{2}-4ac}\div2a\]

OpenStudy (cj49):

a=3 b=-14 c=8 juz plug in d values u ll gt d vale of roots of x

OpenStudy (anonymous):

How do you simplify this, (3x^2-14x+8) / (2x-3)

OpenStudy (anonymous):

yes it's factorable: \[3x^2 -14x+8=> 3x^2 -12x-2x+8 = 3x(x -4) -2(x-4) =>(x-4)(3x-2)\] simplifying it will give: \[\frac{ (x-4)(3x-2) }{2x-3 }\] and thats the limit

OpenStudy (triciaal):

3x^2-14x+8, is this factorable yes = (3 x-2)(x - 4) as above